OpenStudy (anonymous):

Find the positive value for X such that x/2=3/x-5

OpenStudy (shubhamsrg):

cross-multiply, thats your first step

OpenStudy (anonymous):

So you get x squared, I dont know what to do now, beacause its x*x-5

OpenStudy (radar):

Continue on: \[x ^{2}-5x=6\]

OpenStudy (radar):

\[x^{2}-5x-6=0\]

OpenStudy (radar):

Now solve using factoring. select the factor that will produce the positive result.

OpenStudy (radar):

can you factor it?

OpenStudy (anonymous):

no

OpenStudy (radar):

(x-6)(x+1) note that xx=x^2, -6(1)= -6 and that -6 + 1=-5.

OpenStudy (radar):

So now (x-6)(x+1)=0

OpenStudy (anonymous):

huh?

OpenStudy (radar):

Solve for x. the positive value as per request of the problem.

OpenStudy (anonymous):

no the part before that?

OpenStudy (radar):

I was attempting to show you that (x-6)(x+1) were the factors of x^2-5x-6